OpenStudy (anonymous):

verifying an identity (cos x/1+sin x ) + (1+sinx/cosx) = 2 sec x So I started on the left and ended up with 2 + 2 sin x/ 1+ 2 sinx + cos x. I'm not sure what to do next? Can anyone spot me where I'm wrong?

zepdrix (zepdrix):

\[\large \frac{\cos x}{1+\sin x}+\frac{1+\sin x}{\cos x}\]Hmm I'm not really sure what you did to get that result... hmmm

zepdrix (zepdrix):

Oh ok I think I see where you got that :3 Doing something like this I guess?\[\large \frac{\cos^2 x}{\cos x(1+\sin x)}+\frac{(1+\sin x)^2}{\cos x(1+\sin x)}\]

OpenStudy (anonymous):

Here is what got for the top cos^2 +1+2sin x + sin 2 x than 1 - sin^2x + 1 + 2 sin x + 2 sin ^ 2 x

OpenStudy (anonymous):

Yes I Am ^

zepdrix (zepdrix):

Hmm I think the approach we want to take is a little different. We're going to get a common denominator using a different method, it will just kind of, show up. \[\large \frac{\cos x}{1+\sin x}+\frac{1+\sin x}{\cos x}\]Let's multiply the left term (top and bottom) by the conjugate of the bottom.\[\large \color{royalblue}{\frac{1-\sin x}{1-\sin x}}\cdot \frac{\cos x}{1+\sin x}+\frac{1+\sin x}{\cos x}\]

zepdrix (zepdrix):

\[\large \frac{\cos x(1-\sin x)}{1-\sin^2 x}+\frac{1+\sin x}{\cos x}\]Can you see the next step? Hint ~ Important identity involving \(\large 1-\sin^2x\).

OpenStudy (anonymous):

Hmm I'm guessing 1-sin^2x Cancels out on the right side

zepdrix (zepdrix):

Recall the first identity they teach you in Trig.\[\large \cos^2x+\sin^2x=1\]From this identity we can derive this,\[\large 1-\sin^2x=\cos^2x\]So we're going to use this identity to replace our \(\large 1-\sin^2x\) with \(\large \cos^2 x\).

OpenStudy (anonymous):

Oh! Thats right, what occurs next? Excuse my slowness I'm having a tough time with these identities

zepdrix (zepdrix):

\[\large \frac{\cos x(1-\sin x)}{\color{royalblue}{1-\sin^2 x}}+\frac{1+\sin x}{\cos x} \qquad \rightarrow \qquad \frac{\cos x(1-\sin x)}{\color{royalblue}{\cos^2x}}+\frac{1+\sin x}{\cos x}\] We'll cancel out a cosine from the top and bottom of the left term.\[\large \frac{\cancel{\cos x}(1-\sin x)}{\cos^{\cancel21}x}+\frac{1+\sin x}{\cos x} \qquad \rightarrow \qquad \frac{(1-\sin x)}{\cos x}+\frac{1+\sin x}{\cos x}\] How bout that! We were able to get a common denominator by doing something rather sneaky :D

OpenStudy (anonymous):

no way you could do that?! There are hidden rules in these identities... I'm guessing the top sin would cancel out and you would be left with 1/ cos x^2?

OpenStudy (anonymous):

oopsie cos^x

zepdrix (zepdrix):

Yes the sine's on top will cancel, but the 1's will add together, giving us,\[\large \frac{2}{\cos x}\] Yes good catch, no square on the bottom. Just adding.

OpenStudy (anonymous):

alright.. hmm so that would translate to 2 sec x? according to the properties?

zepdrix (zepdrix):

\[\large \frac{(1-\sin x)}{\cos x}+\frac{1+\sin x}{\cos x} \qquad \rightarrow \qquad \frac{1-\sin x+1+\sin x}{\cos x}\]

zepdrix (zepdrix):

\[\large \frac{2}{\cos x} \qquad = \qquad 2 \cdot \frac{1}{\cos x}\]Then we have to remember another important identity.\[\large \sec x=\frac{1}{\cos x}\]

OpenStudy (anonymous):

the 2 comes in front am i right? :3

zepdrix (zepdrix):

Yes :o

OpenStudy (anonymous):

Thank you very much ! :DDD

zepdrix (zepdrix):

c: